# ThinkRec: Thinking-based recommendation via LLM

### v3

@@ 3.3.2 Instance Wise Expert Fusion
$\mathrm{H}(\mathbf{w}^{u})=-\sum_{n=1}^{N}w_{n}^{u}\log{w_{n}^{u}},$ (9)
$\mathrm{LoRA}^{u}=\left\{\begin{array}[]{ll}\mathrm{LoRA}_{global},&\mathrm{averaged}\\ \mathrm{LoRA}_{argmax(\mathbf{w}^{u})},&\mathrm{concentrated}\\ \sum_{n=1}^{N}w_{n}^{u}\mathrm{LoRA}_{n},&\mathrm{otherwise}\end{array}\right..$ (10)
4 Experiments
@@
4.1 Experimental Setup
Datasets.
@@ 4.1 Experimental Setup
Table 1: Statistics of Datasets.
We conduct experiments on three datasets: ML1M refers to MovieLens-1M [^2] https://grouplens.org/datasets/movielens/1m/. Yelp refers to the Yelp Open dataset [^3] https://business.yelp.com/data/resources/open-dataset/. Book refers to the “book” subset of the Amazon Product Review dataset [^4] https://nijianmo.github.io/amazon/index.html. ML1M contains user ratings on movies, collected between 2000 and 2003, with ratings on a scale of 1 to 5. We convert these ratings into binary labels using a threshold of 3. Yelp includes user reviews, ratings for businesses such as restaurants and retail shops, as well as textual information about the businesses. We convert these ratings into binary labels using a threshold of 3. Book compiles user reviews of books from Amazon, collected between 1996 and 2018, with review scores ranging from 1 to 5. We transform these review scores into binary labels using a threshold of 4. Ratings greater than the threshold are labeled as “positive” (y = 1), while the rest are labeled as “negative” (y = 0).
To better simulate real-world recommendation scenarios and prevent data leakage [31], we split the dataset into training, validation, and testing sets based on the interaction timestamp. Specifically, for Amazon-Book, given its large scale, we just preserve interactions from the year 2017, allocating the first 10 months for training, and the remaining 2 months for validation and testing. As for Yelp, we preserve interactions from 2010 to 2022, allocating the first 10 years for training, and the remaining 2 years for validation and testing. Given the sparse nature of Book and Yelp, we filtered out users and items with fewer than 20 interactions. As for ML1M, we preserve the interactions from the most recent twenty months, using the first 10 months for training, the remaining 10 months for validation and testing. The statistical information on the processed datasets is in Table 1. $\sigma$, $\mu$ refer to the standard deviation and mean of user interactions.
Evaluation Metrics. We employ four widely used recommendation metrics: AUC, UAUC, Normalized Discounted Cumulative Gain (NDCG), and Mean Average Precision (MAP). And we employ METEOR [32] and BLEURT [33] to measure the generated reasons. METEOR incorporates synonym matching and word order. BLEURT is fine-tuned on human judgments to directly predict text quality.
Baselines. To fully evaluate the proposed method ThinkRec, we compare it with traditional collaborative methods, and LLM-based methods. We extend LLM-based methods to LLM Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for a fair comparison and tune the LLM with LoRA to manage computational costs.
  • MF [28] is Matrix Factorization, a latent factor-based collaborative filtering method.
  • LightGCN [29] A representative graph-based collaborative filtering method, which uses a simplified graph convolutional neural network to enhance the modeling of user interest.
  • SASRec [34] A representative sequential recommendation method, which uses self-attention to encode sequential patterns to model user interest.
  • Prompt4NR [35] It uses both fixed and soft prompts to utilize traditional Language Models for recommendation.
  • TALLRec [11] This is a state-of-the-art LLMRec method that aligns LLM with recommendations through aligns LLM with recommendations through instruction tuning.
  • CoLLM [5] It effectively integrates collaborative information into LLMs by harnessing the capability of external traditional models to capture the information.
Implementation Details We implement all the compared methods using PyTorch 2.5. We adopt BCE when not otherwise specified. LLM-based methods are optimized using AdamW, while other models use Adam. For hyperparameters of traditional methods, we explore the learning rate within the range of [1e-2, 1e-3, 1e-4], the embedding size within the range of [64, 128, 256], and weight decay within the range of [1e-2, 1e-3, . . . , 1e-7]. As for LLM-based methods, we set the learning rate to 1e-4, and weight decay to 1e-3 to align with CoLLM. For SASRec, we establish the maximum length of historical interaction sequences to 25. We adopt TALLRec’s practice of setting the maximum sequence length to 10 for all other methods. Regarding other specific parameters of the baseline models, we adhere to the configurations outlined in their original papers. For the LoRA module, we follow the same configuration as CoLLM, setting $r,alpha,dropout,target\ modules$ to 8, 16, 0.05, and “[q proj, v proj]”, respectively. We set the sample rate of reasoning data and recommendation data to 0.2 and 0.8, and the weight of each loss function as follows: $\alpha=0.1,\beta=0.9,\eta=0.9,\gamma=0.1$. We set the temperature coefficient $\tau$ to $0.1$. And we set the sampling ratio of reasoning data to 0.2.
@@
Ours 0.7764 0.6775 0.7747 0.4774 0.6955 0.6065 0.8585 0.2826 0.8302 0.5705 0.6858 0.2977
Table 2: Comparison of prediction performance between ThinkRec and the baselines across the three evaluation datasets. The best results are highlighted in bold and sub-optimal results are underlined.
4.2 Overall Performance (RQ1)
In this section, we compare our ThinkRec method with baseline methods on the accuracy of recommendations and the quality of generated reasons.
4.2.1 Accuracy of Recommendation
@@
The results of ThinkRec and SOTA recommendations are in Table 2. We observe some improvement in our method on all three datasets. Notably, ThinkRec improves the previous SOTA CoLLM by +.0582 (9.13%) on AUC in Yelp and +.0623 (8.72%) on AUC in ML1M, demonstrating substantial gains in global ranking accuracy. CoLLM, which integrates collaborative signals into LLMs using external traditional models, achieves the second-best AUC on ML1M (0.7141) and Book (0.7830), confirming its effectiveness in leveraging user–item interaction patterns. However, its performance in user-level ranking metrics is relatively less competitive, especially on ML1M and Yelp, where ThinkRec offers more personalized modeling through instance-wise expert fusion.
@@
Other LLM-based methods, such as Prompt4NR and TALLRec, also demonstrate competitive performance, particularly on datasets with rich textual item metadata such as Book. For instance, TALLRec achieves the highest NDCG@5 (0.7683) on ML1M, and both TALLRec and Prompt4NR slightly outperform ThinkRec in MAP@5 on Book. These results suggest that instruction tuning and prompt engineering are beneficial when item descriptions provide substantial semantic context.
@@
Turning to traditional recommendation models, SASRec clearly outperforms MF and LightGCN, especially on ML1M and Yelp. Its self-attention-based sequential modeling effectively captures temporal patterns in user behavior, yielding the best UAUC (0.6687) and MAP@5 (0.4747) among non-LLM baselines on ML1M. However, its performance drops significantly on Book, where user interactions are more sparse and less sequential, revealing the limitations of sequence-only models.
@@
Overall, ThinkRec delivers the most balanced and robust performance across all datasets and evaluation metrics. It achieves the best performance of almost all metrics on ML1M and Yelp, and also obtains the top AUC and UAUC on Book. Its consistent top-tier results confirm the effectiveness of combining thinking activation and expert fusion. These components jointly enhance both global and user-specific ranking quality, making ThinkRec a scalable and interpretable recommendation.
4.2.2 Quality of Generated Reasons
@@
Ours 0.0333 0.3104 0.0616 0.2683 0.0546 0.2828
Table 3: Evaluation quality of generated reasons. "M" refers to "METEOR" and "B" refers to "BLEURT".
Table 3 summarizes the quality of the reasons generated by the LLM-based recommenders. Compared with the generated reasons from QwQ using the METEOR and BLEURT metrics, the reasons generated by our method significantly outperform those of the three LLM-based baselines. Our method achieves an average relative improvement of 56.54% on METEOR and 23.35% on BLEURT across all datasets, suggesting better fluency, coherence, and semantic relevance.
These results validate the effectiveness of our thinking activation mechanism, which explicitly aligns recommendations with user-centric reasoning via joint training on reasoning-augmented samples. The improvement in both syntactic and learned semantic metrics confirms that ThinkRec not only provides accurate recommendations but also produces more coherent, grounded, and human-aligned explanations, a crucial step toward reasonable and trustworthy LLM-based recommendations.
4.3 In-depth Analysis
4.3.1 Ablation Studies (RQ2)
@@
Ours 0.6775 0.7747 0.4774 0.6065 0.8585 0.2826 0.5705 0.6858 0.2977
Table 4: Ablation studies of key components in ThinkRec. "N" refers to "NDCG", "M" refers to "MAP".
To evaluate the importance of explicit reasoning in recommendation, we ablate the "thinking" component of ThinkRec (w/o think), which disables reasoning supervision.
As shown in Table 4, this leads to significant performance degradation across all datasets. For example, UAUC on the Book dataset drops from 0.5705 to 0.4692. Interestingly, this is even lower than the variant where both thinking and expert mechanisms are removed (w/o both). Without thinking, the recommendation task is effectively reduced to binary classification, where enabling multi-expert modeling may overfit shallow interactions, thus compromising generalization.
Figure 3: The influence of performance with the number of experts on Book (left panel) and Yelp (right panel).
4.3.2 Study on the Fusion of Experts (RQ3)
Analysis of the number of experts. As shown in Figure 3, the number of expert groups increases from 1 to 4, the model exhibits a characteristic ’rise-then-fall’ performance trend, revealing the trade-off between personalization capacity and generalization. In the early stages, fine-tuning LoRA modules within user groups significantly enhances the model’s ability to capture diverse preferences, resulting in notable gains in user-level and Top-N ranking metrics such as UAUC, NDCG@5, and MAP. However, with further partitioning, each subgroup receives fewer training samples, making the model prone to overfitting, thereby degrading ranking performance. Notably, the AUC metric consistently decreases with more experts, reflecting the deterioration of the consistency of global representation with expert specialization, validating the inherent tension between "global consistency" and "local specificity" in recommender systems. These results indicate that more experts do not necessarily equate to better performance; instead, the optimal group number should be adjusted according to user behavior diversity and the frequency of interactions.
Figure 4: The influence of performance on the accuracy of grouping (Cohen’s d of grouped datasets).
Analysis of the grouping features. Under a fixed two-group setting, we further investigate how the choice of user grouping features affects model performance. Specifically, we employ user embeddings generated by MF, LightGCN, and SASRec to construct different grouping strategies. As shown in Figure 4, with the increase of group difference (Cohen’s d), the performance of ThinkRec consistently improves. This trend highlights the importance of semantic decoupling among expert groups. When user preferences across groups exhibit stronger heterogeneity, the LoRA modules assigned to each group can learn more complementary preference representations, thereby enhancing the system’s modeling capacity and global discriminative power. In contrast, if the groups are semantically overlapping, expert models tend to learn redundant representations, and the fused output becomes increasingly similar to that of the global model, limiting performance gains. Therefore, leveraging high-quality user behavior modeling methods as the basis for grouping can amplify divergence across user groups, enabling multi-expert systems to achieve better personalized expressiveness while preserving a global perspective.
5 Conclusion
In this work, we explore how to activate the explicit reasoning capabilities of LLMs for recommendation tasks, aiming to improve both interpretability and accuracy. We highlight crucial challenges in existing LLM-based recommendation approaches, including the absence of reasoning supervision and insufficient modeling of fine-grained user preferences. To this end, we propose the ThinkRec, which incorporates a thinking activation framework containing item augmentation and mixed sampling training, an instance-wise expert fusion mechanism based on latent user features. The results in three real-world datasets confirm that ThinkRec performs better in both the accuracy of recommendations and the quality of explanations. We believe this work lays the foundation for a new paradigm that bridges reasoning and recommendation in LLMs. In the future, we plan to extend this line of research toward thinking-based item generation and review-informed user modeling.
Appendix A Pseudocode of ThinkRec
We briefly introduce ThinkRec in Algorithm 1 and 2. First, as shown in Algorithm 1, items are augmented and reasons are generated. Then global and base experts are trained on according datasets. Second, as shown in Algorithm 2, the participation of each expert is calculated according to their representations and the user’s latent feature. Then, user-specific experts are specified according to participation and thresholds.
1
Algorithm 1 Thinking Activation
@@ Appendix A Pseudocode of ThinkRec
3 Target: Metadata of each item $i_{d}$ $\mapsto$ Description of each item $i_{k}$ Input: Metadata of each item $i_{d}$ Output: Description of each item $i_{k}$
4
[…]
6 Target: Interaction dateset $\mathcal{S}$ $\mapsto$ Reasoning dateset $\mathcal{S}_{r}$ Input: Interaction dateset $\mathcal{S}$ Output: Reasoning dateset $\mathcal{S}_{r}$
7
[…]
9 Target: Backbone model $\mathrm{LLM}(\cdot)$, collaborative encoder $f_{\psi}(\cdot;\mathcal{S})$, dataset $\mathcal{S},\mathcal{S}_{r}$ $\mapsto$ Global expert $\mathrm{LoRA}_{global}$, base experts $\{\mathrm{LoRA}_{1:N}\}$ Input: Backbone model $\mathrm{LLM}(\cdot)$, collaborative encoder $f_{\psi}(\cdot;\mathcal{S})$, dataset $\mathcal{S},\mathcal{S}_{r}$ Output: Global expert $\mathrm{LoRA}_{global}$, base experts $\{\mathrm{LoRA}_{1:N}\}$
10
1
Algorithm 2 Instance-wise Experts Fusion
[…]
3 Target: User embedding $\mathbf{e}_{s}^{u}$, representations of experts $\{\mathcal{E}_{1:N}\}$ $\mapsto$ Particication of experts $\mathbf{w}^{u}$ Input: User embedding $\mathbf{e}_{s}^{u}$, representations of experts $\{\mathcal{E}_{1:N}\}$ Output: Particication of experts $\mathbf{w}^{u}$
4
[…]
6 Target: Particication of experts $\mathbf{w}^{u}$, Global expert $\mathrm{LoRA}_{global}$, base experts $\{\mathrm{LoRA}_{1:N}\}$ $\mapsto$ Instance-wise Expert $\mathrm{LoRA}^{u}$ Input: Particication of experts $\mathbf{w}^{u}$, Global expert $\mathrm{LoRA}_{global}$, base experts $\{\mathrm{LoRA}_{1:N}\}$ Output: Instance-wise Expert $\mathrm{LoRA}^{u}$
7

### v4

@@ 3.3.2. Instance Wise Expert Fusion
(9) $\mathrm{H}(\mathbf{w}^{u})=-\sum_{n=1}^{N}w_{n}^{u}\log{w_{n}^{u}},$
(10) $\mathrm{LoRA}^{u}=\left\{\begin{array}[]{ll}\mathrm{LoRA}_{global},&\mathrm{averaged}\\ \mathrm{LoRA}_{argmax(\mathbf{w}^{u})},&\mathrm{concentrated}\\ \sum_{n=1}^{N}w_{n}^{u}\mathrm{LoRA}_{n},&\mathrm{otherwise}\end{array}\right.$.
3.4. Staged Training
In this section, we describe the training process. First, we train a global LoRA $\mathrm{LoRA}_{global}$ on the full dataset under pure text conditions. In parallel, we train a collaborative model $f_{\psi}(\cdot;\mathcal{S})$ on the full dataset under ID-based conditions. Next, user features extracted from the collaborative model are used to group the data, and the grouped data are employed to fine-tune the last 8 layers of $\mathrm{LoRA}_{global}$, yielding a set of base LoRAs $\{\mathrm{LoRA}_{1:N}\}$. Finally, leveraging the global dataset with injected collaborative signals, we train a projector $proj_{\phi}(\cdot)$ that projects collaborative embeddings into the LLM language space, while keeping $\mathrm{LoRA}_{global}$ fixed.
4. Experiments
@@
4.1. Experimental Setup
4.1.1. Datasets.
We conduct experiments on three datasets: ML1M (MovieLens-1M) [^2] https://grouplens.org/datasets/movielens/1m/, Yelp (Yelp Open Dataset) [^3] https://business.yelp.com/data/resources/open-dataset/, and Book (the “Book” subset of the Amazon Product Review dataset) [^4] https://nijianmo.github.io/amazon/index.html. To better simulate real-world recommendation scenarios and avoid data leakage (Ji et al., 2023b), we split each dataset into training, validation, and testing sets according to interaction timestamps. Specifically, for Amazon-Book, due to its large scale, we retain interactions from 2017 only, using the first 10 months for training and the last 2 months for validation and testing. For Yelp, we retain interactions from 2010 to 2022, assigning the first 10 years to training and the last 2 years to validation and testing. For ML1M, we keep interactions from the most recent 20 months, with the first 10 months for training and the remaining 10 months for validation and testing. To address sparsity in Book and Yelp, we filter out users and items with fewer than 20 interactions. The statistics of the processed datasets are summarized in Table 1, where $\mu$ and $\sigma$ denote the mean and standard deviation of user interactions, respectively.
Regarding labels, ML1M contains user ratings on movies, collected between 2000 and 2003, with ratings on a scale of 1 to 5. We convert these ratings into binary labels using a threshold of 3. Yelp includes user reviews, ratings for businesses such as restaurants and retail shops, as well as textual information about the businesses. We convert these ratings into binary labels using a threshold of 3. Book compiles user reviews of books from Amazon, collected between 1996 and 2018, with review scores ranging from 1 to 5. We transform these review scores into binary labels using a threshold of 4. Ratings greater than the threshold are labeled as “positive” (y = 1), while the rest are labeled as “negative” (y = 0).
Table 1. Statistics of Processed Datasets.
@@ 4.1. Experimental Setup
Table 2. Comparison of prediction performance between ThinkRec and the baselines across the three evaluation datasets. The best results are highlighted in bold and sub-optimal results are underlined.
@@
[…]
4.1.2. Baselines.
To fully evaluate the proposed method ThinkRec, we compare it with traditional methods and LLM-based methods:
  • MF. (Koren et al., 2009) This refers to Matrix Factorization, a representative latent factor-based collaborative filtering method.
  • LightGCN. (He et al., 2020) A representative graph-based collaborative filtering method, which uses a graph convolutional neural network to enhance the modeling of user interest.
  • SASRec. (Kang and McAuley, 2018) A representative sequential recommendation method, which uses self-attention to encode sequential patterns to model user interest.
  • gSASRec. (Petrov and Macdonald, 2023) A graph-enhanced sequential recommendation model that augments SASRec by incorporating item–item transition graphs to capture sequential dependencies.
  • Prompt4NR. (Zhang and Wang, 2023) It uses both fixed and soft prompts to utilize traditional Language Models for recommendation.
  • TALLRec. (Bao et al., 2023) This is a state-of-the-art LLMRec method that aligns LLM with recommendations through aligns LLM with recommendations through instruction tuning.
  • CoLLM. (Zhang et al., 2025a) It effectively integrates collaborative information into LLMs by harnessing the capability of external traditional models to capture the information.
We extend the above LLM-based method to LLM Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for a fair comparison and tune the LLM with LoRA to manage costs.
4.1.3. Evaluation Metrics.
We employ four widely used recommendation metrics: AUC, UAUC, Normalized Discounted Cumulative Gain (NDCG), and Mean Average Precision (MAP). And we employ METEOR (Banerjee and Lavie, 2005) and BLEURT (Sellam et al., 2020) to measure the generated reasons. METEOR incorporates synonym matching and word order. BLEURT is fine-tuned on human judgments to directly predict text quality.
4.1.4. Implementation Details
We implement all the compared methods using PyTorch 2.5. We adopt BCE when not otherwise specified. LLM-based methods are optimized using AdamW, while other models use Adam. For hyperparameters such as learning rate, embedding size, and weight decay of traditional methods, we conduct grid search over commonly used ranges, and follow the original papers for other baseline-specific settings. For LLM-based methods, we adopt CoLLM’s configuration for optimizer, sequence truncation and LoRA settings. Full hyperparameter settings and search ranges are provided in the Appendix D.
4.2. Overall Performance (RQ1)
4.2.1. Accuracy of Recommendation
@@
The results of ThinkRec and SOTA recommendations are in Table 2. We observe improvement in our method on all three datasets. Notably, ThinkRec improves the previous SOTA CoLLM by +.0582 (9.13%) on AUC in Yelp and +.0623 (8.72%) on AUC in ML1M, demonstrating substantial gains in global ranking accuracy. CoLLM, which integrates collaborative signals into LLMs using external traditional models, achieves the second-best AUC on ML1M (0.7141) and Book (0.7830), confirming its effectiveness in leveraging user–item interaction patterns. However, its performance in user-level ranking metrics is relatively less competitive, especially on ML1M and Yelp, where ThinkRec offers more personalized modeling through instance-wise expert fusion.
@@
Other LLM-based methods, such as Prompt4NR and TALLRec, also demonstrate competitive performance, particularly on datasets with rich textual item metadata such as Book. For instance, TALLRec achieves the highest NDCG@5 (0.7683) on ML1M among all methods except ThinkRec, and both TALLRec and Prompt4NR slightly outperform ThinkRec in MAP@5 on Book. These results suggest that instruction tuning and prompt engineering are beneficial when item descriptions provide substantial semantic context.
@@
Turning to traditional recommendation models, SASRec clearly outperforms MF and LightGCN, especially on ML1M and Yelp. Its self-attention-based sequential modeling effectively captures temporal patterns in user behavior, yielding the best UAUC (0.6687) and MAP@5 (0.4747) among non-LLM baselines on ML1M. However, its performance declines significantly on the Book dataset, where user interactions are sparser and exhibit weaker sequential patterns, revealing the inherent limitations of models that rely solely on sequential behavior modeling.
@@
Overall, ThinkRec delivers the most balanced and robust performance across all datasets and evaluation metrics. It achieves the best performance of almost all metrics on ML1M and Yelp, and also obtains the top AUC and UAUC on Book. Its consistent top-tier results confirm the effectiveness of combining thinking activation and instance-wise expert fusion. These components jointly enhance both global and user-specific ranking quality, making ThinkRec a scalable and interpretable recommendation.
4.2.2. Quality of Generated Reasons
Table 3 summarizes the quality of the reasons generated by the LLM-based recommenders. Compared with the generated reasons from QwQ using the METEOR and BLEURT metrics, the reasons generated by our method significantly outperform those of the three LLM-based baselines. Our method achieves an average relative improvement of 56.54% on METEOR and 23.35% on BLEURT across all datasets, suggesting better fluency, coherence, and semantic relevance. These results validate the effectiveness of our thinking activation mechanism, which explicitly aligns recommendations with user-centric reasoning via joint training on reasoning-augmented samples. The improvement in both syntactic and semantic metrics confirms that ThinkRec not only provides accurate recommendations but also produces more coherent, grounded, and human-aligned explanations, a crucial step toward reasonable and trustworthy LLM-based recommendations.
Table 3. Quality evaluation of generated reasons. ”M” refers to ”METEOR” and ”B” refers to ”BLEURT”.
@@ 4.2.2. Quality of Generated Reasons
[…]
Figure 3. The influence of performance with the number of experts on Book (left panel) and Yelp (right panel).
The influence of performance with the number of experts on Book (left panel) and Yelp (right panel).
4.2.3. Case Study
Among existing LLM-based recommendations, CoLLM and Prompt4NR yield disorganized symbols. TALLRec frequently generates sentences with unrelated elements, such as code or hallucinated facts, failing to reflect coherent reasoning. In contrast, ThinkRec demonstrates structured, step-by-step reasoning aligned with user history and target item semantics, enabling it to produce accurate and interpretable recommendations. A real example is listed in the Appendix C.
4.3. In-depth Analysis
4.3.1. Ablation Studies (RQ2)
To evaluate the importance of explicit reasoning in recommendation, we ablate the ”thinking” component of ThinkRec, which disables reasoning supervision (w/o think). As shown in Table 4, this leads to significant performance degradation in all datasets. For example, UAUC on the Book dataset drops from 0.5705 to 0.4692. Interestingly, this is even lower than the variant in which both thinking and expert mechanisms are removed (w/o both). This counterintuitive result suggests that, in the absence of thinking, the recommendation task degenerates into binary classification, where multi-expert modeling risks overfitting superficial interactions and thereby undermining generalization.
Table 4. Ablation studies of key components in ThinkRec. ”N” refers to ”NDCG”, ”M” refers to ”MAP”.
@@ 4.3.1. Ablation Studies (RQ2)
Table 5. Metrics before and after model parameter updates.
Models UAUC NDGC@5 MAP@5
global 0.5098 0.6462 0.2657
global* 0.5119 (0.41%) 0.6542 (1.24%) 0.2704 (1.77%)
2 experts 0.5380 0.6641 0.2805
3 experts 0.5669 (5.37%) 0.6890 (3.75%) 0.2950 (5.17%)
4.3.2. Study on the Fusion of Experts (RQ3)
Necessity of multiple experts.
As shown in Figure 3, ThinkRec achieves clear improvements at the optimal number of expert groups (2–3), particularly in UAUC. Prior works(Guo et al., 2017; Zhou et al., 2018) emphasize UAUC as a key performance indicator, where even small gains are considered valuable in practice. In real-world scenarios. Moreover, the multi-expert design improves robustness by capturing diverse user behaviors and enables modular expansion without retraining the full model. We divided the Book dataset into three groups: two groups of original data and one group of newly introduced data. We then tested the performance of the unified and multi-expert models separately. As shown in Table 5, updating only the unified model parameters provides minimal improvement as the number of users increases. In contrast, the multi-expert model has higher original metrics and shows greater improvement.
Analysis of the number of experts.
As shown in Figure 3, the number of expert groups increases from 1 to 4, the model exhibits a characteristic ’rise-then-fall’ performance trend, revealing the trade-off between personalization capacity and generalization. In the early stages, fine-tuning LoRA modules within user groups significantly enhances the model’s ability to capture diverse preferences, resulting in notable gains in user-level and Top-N ranking metrics such as UAUC, NDCG@5, and MAP. However, with further partitioning, each subgroup receives fewer training samples, making the model prone to overfitting, thereby degrading ranking performance. Notably, the AUC metric consistently decreases with more experts, reflecting the deterioration of the consistency of global representation with expert specialization, validating the inherent tension between ”global consistency” and ”local specificity” in recommender systems. These results indicate that more experts do not necessarily equate to better performance; instead, the optimal group number should be dynamically adjusted according to user behavior diversity, and the frequency of interactions.
Figure 4. The influence of performance on the accuracy of grouping (Cohen’s d of grouped datasets).
The influence of performance on the accuracy of grouping (Cohen’s d of grouped datasets).
Analysis of the grouping features.
Under a fixed two-group setting, we further investigate how the choice of user grouping features affects model performance. We employ user embeddings generated by MF, LightGCN, and SASRec to construct different grouping strategies. As shown in Figure 4, with the increase of group difference (Cohen’s d), the performance of ThinkRec consistently improves. This trend highlights the importance of semantic decoupling among expert groups. When user preferences across groups exhibit stronger heterogeneity, the LoRA modules assigned to each group can learn more complementary preference representations, thereby enhancing the system’s modeling capacity and global discriminative power. Therefore, leveraging high-quality user behavior modeling methods as the basis for grouping can amplify divergence across groups, enabling multi-expert systems to achieve better personalized expressiveness while preserving a global perspective.
4.3.3. Cross‑Domain Generalization
In practical recommender systems, distinct models are often deployed for different domains due to significant variations in user behavior across scenarios. In multi-domain settings, a single model is typically trained on a mixture of item types to handle all cases. In contrast, ThinkRec introduces a more nuanced strategy: each base expert is trained on data grouped by user latent features that are specific to the target domain, allowing the model to capture fine-grained, domain-specific preference patterns while maintaining a unified reasoning backbone. To rigorously evaluate cross-domain generalization, we conducted experiments using the global expert. Specifically, we test the metrics in the datasets themselves and use experts trained on the ML1M dataset to test the metrics on the Book and Yelp datasets, as shown in Table 6. The results demonstrate that the reasoning paths learned from ML1M transfer effectively to other domains, retaining more than $90\%$ of the performance achieved with domain-specific training. This demonstrates ThinkRec’s strong generalization capacity and its ability to leverage shared reasoning structures across diverse web environments, reducing the need for domain-specific retraining while maintaining robust recommendation quality.
Table 6. Cross‑Domain Generalization Evaluation.
Datasets ML1M Yelp ML1M-¿Yelp Book ML1M-¿Book
AUC 0.7843 0.7104 0.6709 (94.4%) 0.8369 0.8001 (95.6%)
NDGC@5 0.7740 0.8562 0.7717 (90.1%) 0.6801 0.6197 (91.1%)
METEOR 0.0970 0.0892 0.0876 (98.2%) 0.1028 0.1040 (101.2%)
BLEURT 0.4043 0.3491 0.3421 (98.0%) 0.3489 0.3562 (102.1%)
4.3.4. Inference Efficiency
Although ThinkRec leverages multiple LoRA‑tuned experts during training, inference requires only a single LLM forward pass with the fused LoRA adapter (via weighted sum of a few small matrices), and a lightweight gating computation over user embeddings. Because LoRA adapters add only a small fraction ($1.5\%$) of extra parameters and incur negligible runtime overhead, the overall latency increase is minimal (empirically $10\%$ relative to a standard LoRA‑tuned LLM) while delivering substantially richer, interpretable outputs.
5. Conclusion
The core contribution of this research is advancing LLM-based recommendation systems from an intuitive, matching-based ”System 1” paradigm to a ”System 2” paradigm characterized by explicit reasoning. Through our proposed thinking activation mechanism and instance-wise expert fusion strategy, ThinkRec deeply couples the recommendation task with the reasoning capabilities of language models. Experimental results confirm that this approach not only significantly improves recommendation accuracy but, crucially, enhances system transparency and user trust by generating traceable reasons. We believe that future research will move towards deeper cognitive insights or exploring training-free mechanisms like activation steering for personalized reasoning(Jinghao et al., 2025), enabling systems to perform more rigorous logical reasoning. Moreover, based on the principles of automatic and personalized selection, deeper levels of Collective Intelligence, such as in Multi-Agent Systems, can be explored. ThinkRec builds a solid bridge for this leap. We provide the ethical use of data in Appendix A.
Acknowledgements.
This work was supported by National Natural Science Foundation of China (No. 62402429, U24A20326, 62441236), the Key Research and Development Program of Zhejiang Province (No. 2025C01026, 2024C03270), Ningbo Yongjiang Talent Introduction Programme (2023A-397-G), Young Elite Scientists Sponsorship Program by CAST (2024QNRC001), and partially supported by MYbank, Ant Group. The author gratefully acknowledges the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
Appendix A Ethical Considerations
We use only publicly available benchmark datasets (e.g., ML1M, Yelp, Book) that are widely adopted in the recommendation community. These datasets do not contain any personally identifiable information (PII) or sensitive attributes, and our study does not involve human participants or subjects, thus not requiring informed consent. All data usage strictly complies with ACM’s Publications Policy on Research Involving Human Participants and Subjects.